# Random-Set Neural Networks (RS-NN)

### v2

@@ 4.1 Implementation
RS-NN $\mathbf{79.92}$ Pignistic entropy $3.153\pm 2.549$ $\mathbf{0.1416}$ $\mathbf{60.38}$ $\mathbf{55.16}$ $3.911\pm 2.318$
@@
ENN $71.82$ Mean Entropy $\mathbf{1.395}\pm\mathbf{1.510}$ $0.5961$ $54.67$ $43.73$ $1.617\pm 1.597$
@@
CNN $98.90\pm 0.04$ $92.08\pm 0.42$ $90.89\pm 0.10$ $65.50\pm 0.08$ $78.56$ $94.34$ $\mathbf{1.91}\pm\mathbf{0.03}$
R-BNN https://doi.org/10.48550/arxiv.1506.02557 $90.32\pm 0.14$ $75.21\pm 0.13$ $82.12\pm 0.08$ $50.45\pm 0.02$ $62.17\pm 0.76$
LB-BNN hobbhahn2022fast $99.58\pm 0.04$ $89.95\pm 0.81$ $90.49\pm 0.42$ $59.89\pm 1.96$ $72.48$ $90.85$ $7.11\pm 0.89$
ENN osband2024epistemic $99.07\pm 0.11$ $91.55\pm 0.60$ $91.49\pm 0.19$ $68.02\pm 0.26$ $71.82$ $89.48$ $3.10\pm 0.03$
RS-NN $\mathbf{99.71}\pm\mathbf{0.03}$ $\mathbf{93.53}\pm\mathbf{0.09}$ $\mathbf{94.22}\pm\mathbf{0.03}$ $\mathbf{71.61}\pm\mathbf{0.07}$ $\mathbf{79.92}$ $\mathbf{94.47}$ $\mathbf{1.91}\pm\mathbf{0.02}$
@@ 4.2 Comparison with the state-of-the-art on accuracy
Tab. 2 shows RS-NN outperform state-of-the-art Bayesian (R-BNN, LB-BNN) and Ensemble (ENN) methods in terms of test accuracy (%) across all datasets. On ImageNet, RS-NN’s accuracy is remarkably close to CNN (standard ResNet50), highlighting its efficacy on large and complex datasets. It is important to note that the comparison with standard CNNs is not based on their role as uncertainty methods, but as benchmarks, underscoring RS-NN’s effectiveness in achieving high accuracy compared to other uncertainty baselines. Test accuracy for RS-NN is calculated by determining the class with the highest pignistic probability for each prediction and comparing it with the true class. As RS-NN leverages a random set framework which does not require prior assumptions and is more data driven, it can better adapt and capture the inherent complexity of the data, contributing to its superior performance in test accuracy across various datasets (Tab. 2). The pignistic predictions derived from the predicted belief functions appear more reliable and accurate than standard predicted probabilities of a CNN. R-BNN results for ImageNet are not reported in Tab. 2 as evaluating it on large-scale datasets is impractical due to its complexity. Experiments on inference time (ms) per sample over 5 runs (a single forward propagation) on CIFAR-10 dataset demonstrates the fastest inference time (ms) for RS-NN among all models (Tab. 2).
@@ 4.3 Out-of-distribution (OoD) detection
We evaluate out-of-distribution (OoD) detection (Tab. 1) against baselines using AUROC and AUPRC scores. OoD detection identifies data points deviating from the in-distribution (iD) training data by measuring true and false positive rates (AUROC) and precision and recall trade-offs (AUPRC), providing insights into the model’s performance at varying confidence levels (detailed in §C.3).
RS-NN greatly outperforms Bayesian, Ensemble and standard CNN models in OoD detection (Tab. 1), with significantly higher AUROC and AUPRC scores for all iD vs OoD datasets, especially on CIFAR-10 vs SVHN/Intel Image. This demonstrates its effectiveness in identifying data points deviating from training data. In real-world scenarios, encountering OoD instances is inevitable. Reliable OoD detection is crucial for safety-critical applications to prevent erroneous critical decisions on unfamiliar or anomalous data. Recognising unfamiliar data signals the model about situations beyond its training, allowing it to acknowledge its own limitations and enhance its awareness of ignorance, achieved effectively by RS-NN. Tab. 1 also reports the Expected Calibration Error (ECE) (detailed in §C.2). A low ECE indicates that the model’s confidence scores align closely with the actual likelihood of events. RS-NN exhibits the lowest ECE indicating well-calibrated probabilistic predictions.
@@ 4.4 Uncertainty estimation
Entropy. Tab. 1 shows the mean and variance of the entropy distributions for both iD and OoD datasets. Lower entropy values for iD datasets indicate a confident prediction as the model is trained on familiar data, while higher entropy for OoD datasets reflects the model’s uncertainty with unseen data. RS-NN exhibits both low entropy for iD datasets and high entropy for OoD ones. The percentage mean iD/OoD shift in entropy is most pronounced in RS-NN, especially evident in ImageNet, where CNN exhibits high entropy for iD and OoD datasets. RS-NN maintains a desirable iD vs OoD entropy ratio. This ideal behavior, along with superior OoD detection (Sec. 4.3), is a good indication that RS-NN is better at estimating uncertainty induced by domain shift (see §D.5.2).
Credal set width. Fig. 6 displays the estimated credal set width, the difference between the lower and upper probability bounds in (9), for correctly and incorrectly classified classifications of 1000 CIFAR-10 test samples. Incorrect predictions (orange) correlate with larger credal set widths, indicating higher epistemic uncertainty, and are dispersed in the middle of the plot. Correct predictions (blue) concentrate at the top, exhibiting smaller credal intervals as expected. Some instances of correct classification still display a sizable credal set width, signifying elevated uncertainty and low confidence. Tab. 3 reports the credal set widths of RS-NN predictions for iD vs OoD datasets. Larger intervals can be observed for OoD datasets. Note that credal set width is not directly comparable to the entropy or variance of other models, as such metrics have distinct semantics.
@@
Figure 6: Width of credal predictions for 1,000 CIFAR-10 test samples (correctly classified, blue; incorrectly classified, orange).
@@
4.5 Adaptability to large-scale architectures
Tab. 4 shows the adaptability of RS-NN to larger model architectures and its ability to leverage transfer learning. RS-NN demonstrates higher accuracy compared to standard CNNs across various large-scale model architectures, including WideResNet28-10 (WRN-28-10), VGG16, InceptionV3 (IncV3), EfficientNetB2 (ENetB2), and to Vision Transformer (ViT-Base), showcasing its versatility and ease in adopting different architectures and the generality of the random-set concept. Notably, using a pre-trained ResNet50 (Pre-trained R50) model with ImageNet weights, RS-NN achieves higher accuracy on CIFAR-10 compared to using only the architecture (no pre-trained weights).
@@ 4.6 Limitations
The budgeting step can be time consuming but, for a given dataset, it is a one-time procedure prior to training. For large datasets, t-SNE and GMM can be applied to representative samples rather than the entire dataset, as the purpose is merely to identify the most relevant sets, reducing computational demands. To further optimise the method, instead of t-SNE, budgeting can utilize various methods such as autoencoders ghasedi2017deep ; guo2017deep or PCA abdi2010principal for dimensionality reduction. However, despite the initial time investment, the efficiency gained from having a budget applicable to any training scenario is substantial. Another limitation is manually setting the number of focal elements $K$; a dynamic strategy adjusting $K$ based on overlap would enhance flexibility and effectiveness. The budgeting procedure can also be adapted to work with multiple sources of data or a stream of data (§D.4). Alternative methods such as using sparse mass functions itkina2020evidential ; chen2021evidential , could provide a quantitative basis for determining the subsets $A_{K}$ to consider.
@@ 5 Conclusion
This paper proposes a novel Random-Set Neural Network (RS-NN) for uncertainty estimation in classification using belief functions and random sets. Random-set representation is a foundational method that acts as a versatile wrapper and extends beyond CNNs to any model architecture and classification task (e.g, text). To our knowledge, this concept, along with budgeting for optimal selection of sets, is unprecedented. RS-NN outperforms state-of-the-art uncertainty estimation models and the standard CNN in performance and out-of-distribution OoD tests, and scales seamlessly to large-scale architectures and datasets. Given such results, our approach exhibits significant potential impact for safety-critical applications such as medical diagnostics and autonomous driving, where uncertainty estimation and OoD detection is crucial. In this paper, we discuss target-level representation of RS-NNs. Future work involves parameter-level representation of RS-NN and the extension of this concept from classification to detection tasks.
Acknowledgement
This work has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 964505 (E-pi).
Appendix A Further discussion
A.1 Rationale and applicability
This paper aims to establish a foundational approach for integrating belief functions into standard deep learning models, specifically CNNs. We chose to closely adhere to a standard CNN framework to ensure the model’s applicability to large-scale, real-world datasets—a capability often lacking in other uncertainty models. As a result, we achieve performance comparable to that of a standard CNN while providing improved and reliable uncertainty estimation.
However, the proposed method can be thought of as a generalisation of traditional classification. There is nothing limiting it to CNNs. In fact, Tab. 4 shows that it can be easily adapted to work with a wide range of network architectures, including Vision Transformers (ViTs). Similarly, it can also work with simple Multi-Layer Perceptrons (MLPs).
A.2 Statistical guarantees
@@
To complement the extensive discussion in the main paper, we provide here an in-depth discussion on the statistical guarantees underlying RS-NN and the rationale behind our choices for ground-truth representation $\mathbf{bel}$, loss function and training, under how RS-NN learns sets of outcomes.
Unlike, e.g., conformal learning, our RS-NN approach does not, in its current state, provide statistical guarantees such as the coverage of the prediction intervals. However, we can see at least three ways in which this could be done.

### v3

@@ 4.1 Implementation
RS-NN $\mathbf{79.92}$ Pignistic entropy $2.972\pm 2.108$ $\mathbf{0.1416}$ $\mathbf{60.38}$ $\mathbf{55.16}$ $3.659\pm 3.771$
@@
FSVI $62.56$ Predictive Entropy $\mathbf{1.328}\pm\mathbf{1.966}$ $0.3890$ $50.55$ $49.88$ $1.637\pm 1.328$
DE $78.77$ Mean Entropy $1.532\pm 1.325$ $0.1940$ $55.37$ $53.20$ $1.775\pm 1.343$
ENN $71.82$ Mean Entropy $1.395\pm 1.510$ $0.5961$ $54.67$ $43.73$ $1.617\pm 1.597$
@@ 4.2 Comparison with the state-of-the-art on accuracy
Tab. 1 shows that RS-NN outperforms state-of-the-art Bayesian (LB-BNN, FSVI) and Ensemble (DE, ENN) methods in terms of test accuracy (%) across all datasets. Experiments on inference time (ms) per sample, conducted over 5 runs (a single forward propagation) on the CIFAR-10 dataset, show that RS-NN and CNN have the fastest inference times among all models (Tab. 1). Albeit CNN and DE are close to RS-NN in performance (e.g., on ImageNet), CNN tends to be overconfident in incorrect predictions (§E.2), while DE has significantly longer training (Tab. 6, §D) and inference times (Tab. 1). The pignistic predictions derived from predicted belief functions are more reliable and accurate than predicted softmax probabilities of CNN and DE. It is important to note that the comparison with standard CNN is not based on its role as an uncertainty method, but as a benchmark, underscoring RS-NN’s effectiveness in achieving high accuracy compared to other uncertainty baselines. Test accuracy for RS-NN is calculated by determining the class with the highest pignistic probability for each prediction and comparing it with the true class. This proves that, by modelling the epistemic uncertainty about the prediction, we take better into account possible distribution shifts at test time - as a result, the central prediction of the set of probabilities (credal set) associated with the predicted belief function is more likely to be closer to the ground truth. Note that the results in the FSVI paper (Rudner et al., 2022) are based on ResNet18. Although using a pre-trained ResNet50 could enhance FSVI’s performance, it would provide an unfair advantage over the other baseline models in our paper, which were all trained from scratch.
@@ 4.3 Out-of-distribution (OoD) detection
We evaluate our out-of-distribution (OoD) detection (Tab. 2) against baselines using AUROC and AUPRC scores. OoD detection identifies data points deviating from the in-distribution (iD) training data by measuring true and false positive rates (AUROC) and precision and recall trade-offs (AUPRC), indicating the model’s ability to handle unfamiliar data (further detailed in §C.3). As shown in Tab. 2, RS-NN greatly outperforms Bayesian, Ensemble and standard CNN models in OoD detection, with significantly higher AUROC and AUPRC scores for all iD vs OoD datasets, especially on difficult datasets like ImageNet and ImageNet-O (see §E.1). Even high-accuracy models like CNN and DE struggle to differentiate between ImageNet and ImageNet-O. While DE shows comparable AUROC scores on CIFAR-10 vs. SVHN and MNIST vs. F-MNIST/K-MNIST to RS-NN, it has lower AUPRC. While AUROC assesses a model’s ability to distinguish between iD and OoD samples, AUPRC measures both the precision of correctly identified OoD samples and the recall of detected OoD samples, making it a more informative metric. Tab. 2 also reports the various models’ Expected Calibration Error (ECE) (detailed in §C.2). A low ECE indicates that the model’s confidence scores align closely with the actual likelihood of events. RS-NN exhibits the lowest ECE indicating well-calibrated probabilistic predictions.
@@ 4.4 Uncertainty estimation
Entropy. Tab. 2 shows the mean and variance of the entropy distributions for both iD and OoD datasets. Lower entropy values for iD datasets indicate a confident prediction as the model is trained on familiar data, while higher entropy for OoD datasets reflects the model’s uncertainty with unseen data. RS-NN exhibits both low entropy for iD datasets and high entropy for OoD ones. The percentage mean iD/OoD shift in entropy is most pronounced in RS-NN, especially evident in CIFAR-10 vs SVHN/Intel Image and MNIST vs F-MNIST/K-MNIST. For ImageNet vs ImageNet-O, CNN exhibits high entropy for both iD and OoD datasets while RS-NN maintains a desirable iD vs OoD entropy ratio. In fact, Fig. 18, §E.5.1 show that RS-NN has the best iD vs OoD entropy ratio for all datasets. This ideal behavior, along with superior OoD detection (Sec. 4.3), is a good indication that RS-NN is better at estimating uncertainty induced by domain shift (see §E.5.1).
Credal set width. Fig. 6 shows a density plot of estimated credal set widths, the difference between the lower and upper probability bounds in Eq. 9, for correctly and incorrectly classified samples of CIFAR-10 test data. Incorrect predictions (red) correlate with larger credal set widths, indicating higher epistemic uncertainty, and are more dispersed in the plot. Correct predictions (blue) concentrate around smaller values, exhibiting smaller credal intervals as expected. Tab. 3 reports the credal set widths of RS-NN predictions for iD vs OoD datasets. Larger intervals can be observed for OoD datasets. The credal set widths for ImageNet vs ImageNet-O are not as distinct given the nature of these datasets, as detailed in §E.1. Note that credal set width is not directly comparable to the entropy, variance or mutual information of other models, as such metrics have distinct semantics.
@@
Figure 6: Width of credal predictions for CIFAR-10 test data (correctly classified, blue; incorrectly classified, red).
@@
4.5 Scalability to large-scale architectures
Tab. 4 shows the scalability of RS-NN to larger model architectures and its ability to leverage transfer learning. RS-NN outperforms standard CNNs across various large-scale model architectures, including WideResNet28-10 (WRN-28-10), VGG16, InceptionV3 (IncV3), EfficientNetB2 (ENetB2), and Vision Transformer (ViT-Base-16), highlighting its versatility and ease in adopting different architectures, and the generality of the random-set concept. Notably, using a pre-trained ResNet50 (Pre-trained R50) model with ImageNet weights, RS-NN achieves higher accuracy on CIFAR-10 compared to using only the architecture (no pre-trained weights).
@@ 4.6 Limitations
The budgeting step can be time consuming but, for a given dataset, it is a one-time procedure prior to training. For large datasets, t-SNE and GMM can be applied to representative samples rather than the entire dataset, as the purpose is merely to identify the most relevant sets, dramatically reducing computational demands. To further optimise the method, instead of t-SNE, budgeting can utilize various methods such as autoencoders (Ghasedi Dizaji et al., 2017; Guo et al., 2017b) or PCA (Abdi & Williams, 2010) for dimensionality reduction. However, despite the initial time investment, the efficiency gained from having a budget applicable to any training scenario is substantial. Another limitation is manually setting the number of focal elements $K$; a dynamic strategy adjusting $K$ based on overlap would enhance flexibility and effectiveness. The budgeting procedure can also be adapted to work with multiple sources of data or a stream of data (§E.8). Alternative methods such as using sparse mass functions (Itkina et al., 2020; Chen et al., 2021), could provide a quantitative basis for determining the subsets $A_{K}$ to consider.
@@ 5 Conclusion
This paper proposes a novel Random-Set Neural Network (RS-NN) for uncertainty estimation in classification predicting belief functions. Our random-set representation is a foundational approach that acts as a versatile wrapper, applicable to any model architecture and classification task (e.g, text). To our knowledge, this concept, along with budgeting for optimal selection of sets, is unprecedented. RS-NN outperforms state-of-the-art uncertainty estimation models and the standard CNN in performance and out-of-distribution OoD tests, and scales seamlessly to large-scale architectures and datasets. Given such results, our approach exhibits significant potential impact for safety-critical applications such as medical diagnostics and autonomous driving, where uncertainty estimation and OoD detection is crucial. Future work will explore alternative methods for ensuring the validity of belief functions without mass regularisation terms, such as integrating Semantic Probabilistic Layers (Ahmed et al., 2022) to enforce logical constraints in probabilistic predictions, or applying softmax over the computed masses to guarantee non-negativity (see §E.6). A parameter-level representation of RS-NN and the extension of this concept to regression will also be explored.
Reproducibility Statement
To ensure reproducibility, we provide detailed information on the datasets, baseline models, and the training procedure for RS-NN and all baselines used in our experiments in Sec. 4.1, under the subheadings ‘Datasets’ and ‘Baselines, backbones, and training details’ respectively. Additional hyperparameters for RS-NN are provided under ‘Training the RS-NN’ (Sec. 4.1), and for baseline models in §D. For convenience, we summarize the training hyperparameters for all models in Tab. 5. Additionally, the training times for each model are included in Tab. 6D). The code for training RS-NN (train.ipynb), including pre-trained models (for faster evaluation), evaluation script (eval.ipynb), and configuration files are provided as supplementary materials. We also provide a .yml file to set up the environment to run the experiments.
Appendix: Random-Set Neural Networks
This appendix provides a comprehensive overview of Random-Set Neural Networks (RS-NN), including theoretical underpinnings (§A), related work (§B), algorithms (§C), implementation details (§D), additional experimental results (§E) which addresses critical questions to elucidate the functionality and robustness of RS-NN, and an application to text classification (§F).
Appendix A Theoretical explanation of RS-NN
In this section, we provide an overview of belief functions (§A.1), describe the RS-NN learning mechanism (§A.2), and show the relationship between pignistic and credal predictions (§A.3), and provide statistical guarantees for RS-NN by applying conformal prediction to the predicted pignistic probabilities (§A.4).
A.1 Classical probabilities vs Belief functions
In this section, we explain the difference between classical probability measures and non-additive measures, such as random sets. Random sets (Matheron, 1975; Kendall, 1974; Nguyen, 1978; Molchanov, 2005) are non-additive measures (i.e., the additivity property does not hold). Their additional degrees of freedom allow to express the ‘epistemic’ uncertainty about probability values themselves. In fact, an entire field on ‘imprecise probabilities’ (Augustin et al., 2014; Walley, 2000) exist which deals with these kinds of measures.
In a standard classification task, where each input is classified into one of several mutually exclusive classes, the classical probability framework does assume that each class is distinct and independent. The output of a softmax model provides the probabilities for each class, and these probabilities sum to 1. This approach works well when you are confident about the classification, for you have sufficient evidence to assign a clear probability to each class. However, it does not capture the uncertainty or ambiguity that exists when the evidence (as provided by the training set) is insufficient (e.g., because you did not sample similar points at training time, or because the test data is affected by distribution shift).
In Sec. 1 of the paper, we highlight that relying solely on confidence measures (like softmax probabilities) is insufficient for a comprehensive analysis of model predictions. Fig. 2 shows that standard neural networks are often overconfident (Nguyen et al., 2015) even for misclassifications. Our approach uses random sets (which, in the finite case, assume the name of belief functions), as detailed in Sec. 2, to offer a robust framework for assessing prediction reliability.
[…]
Unlike classical probability theory, belief functions operate in a higher space than probability vectors, i.e, it operates in a framework that generalizes classical probability theory. Suppose we know with 100% certainty that an event is either A or B. In a classical probability setting, we might say P(A) = 0.5 and P(B) = 0.5, or we might assign different probabilities to A and B if we have some reason to favor one over the other. With belief functions, we can represent the same scenario differently. We could assign a belief value of 1 to the set {A, B} and 0 to both A and B individually (Bel({A}) = Bel({B}) = 0, Bel({A, B}) = 1). This means that while we know the outcome must be either A or B, we are entirely uncertain about which one it is. Alternatively, we could have Bel({A}) = 0.5, Bel({B}) = 0.5, Bel({A, B}) = 1, reflecting a different state of knowledge or evidence, where we’re somewhat more informed but still not fully certain. In the behavioural interpretation of probability (Walley, 2000), the belief value of an event A is the upper bound to the price one is willing to pay for betting on an outcome A.
Consider the following example (Cuzzolin, 2020): Suppose there is a murder, and three people—Peter, John, and Mary—are suspects. Our hypothesis space is therefore $\Theta$ = {Peter, John, Mary}. A witness testifies that the person he saw was a man, supporting the proposition A = {Peter, John} $\subseteq\Theta$. However, the witness was tested, and the machine reported a 20% chance that he was drunk when he gave his testimony. As a result, we assign 80% belief to the proposition A, representing our uncertainty about whether any of the two could be the murderer.
In classical probability, Kolmogorov’s additive probability theory (Kolmogorov & Bharucha-Reid, 2018) forces us to specify support for individual outcomes, i.e. to distribute this 80% probability between Peter and John individually, even when the evidence (our data) supports set propositions. This is unreasonable – an artificial constraint due to a mathematical model that is not general enough. In the example, we have do not have enough evidence to assign this 80% probability to either Peter or John, nor information on how to distribute it amongst them. The cause is the additivity constraint that probability measures are subject to. This constraint reflects a broader limitation of measure-theoretical probability, which, while effective in many scenarios, struggles with second-order uncertainties and incomplete information. Therefore, we have several other methods to estimate uncertainty, such as the Bayesian method, or using degrees of belief, such as Evidential and belief function methods (see Sensoy et al. (2018), Sec. 3).
This means that belief functions allow us to model situations where we have uncertainty not just about which class an input belongs to, but also about whether we have enough evidence to distinguish between certain classes. This added flexibility is what makes belief functions suitable for single-label classification tasks.
A.2 RS-NN Learning Mechanism
@@
Here we wish to expand on the rationale behind our choices for ground-truth representation $\mathbf{bel}$, loss function and training, and the way in which RS-NN learns sets of outcomes.
Recall from Sec. 3.1 that the ground-truth for a RS-NN is the belief-encoded vector $\mathbf{bel}=\{Bel(A),A\in\mathbb{P}(\mathcal{C})\}$, where $Bel(A)$ is the belief function of a focal set $A$ in the power set $\mathbb{P}$ of classes $\mathcal{C}$. In our method, $Bel(A)$ is 1 if the true class is in subset $A$ and 0 otherwise. Consequently, the belief-encoded ground truth $\mathbf{bel}$ will include multiple occurrences of 1. For instance, in a digit-classification task such as MNIST, if $\{3\}$ is the true class, the belief-encoded ground truth would contain 1s in all sets where $\{3\}$ is present, such as {1, 3}, {0, 3}, {1, 2, 3}, and so forth, 0 for sets not containing $\{3\}$ at all, such as {0, 1}, {1, 2}, {0, 1, 2}, etc.
A.3 Pignistic probability and credal prediction
Figure 7: Upper and lower bounds (Eq. 9), in dotted red, to the probability of the predicted most likely class according to the pignistic prediction (in solid red) for 100 samples of CIFAR-10.
As they have belief values as lower bounds, credal sets induced by belief functions have a peculiar shape (see Fig. 4, right, for a case with 3 classes). Their vertices are induced by the permutations of the elements of the sample space (for us, the set of classes) (Chateauneuf & Jaffray, 1989; Cuzzolin, 2008). Given one such permutation, e.g., $(c_{2},c_{4},c_{1},c_{3})$ for a set of 4 classes, the corresponding probability vector (vertex of the credal prediction) assigns to each class the mass of all the focal sets containing it, but not containing any class preceding it in the permutation order (Wallner, 2005).
For additional illustration, the lower and upper bounds (Eq. 9) to the probability of the top predicted class are plotted in Fig. 7, together with the pignistic probability, for 100 samples of CIFAR-10. The bounds Eq. 9 can be efficiently computed from the finite number of vertices of the credal prediction ((Cuzzolin, 2008), Fig. 4).
A.4 Statistical guarantees
To complement the extensive discussion in the main paper, we provide here an in-depth discussion on how statistical guarantees can be provided in an RS-NN framework.